OpenStudy (anonymous):

For which value of k does the matrix A=[4 k; 2 −2] have one real eigenvalue of algebraic multiplicity 2? let lambda = x x^2 -2x + (-8-2k) = 0

OpenStudy (perl):

The eigenvalues of A are x, where det(A-x*Identity) = 0. $$ \Large{\det \begin{bmatrix} 4-x & k \\2& -2-x \end{bmatrix}=0 \\ ~\\\iff \\~\\ x^2-2x-2k-8=0 } $$

OpenStudy (perl):

now use the quadratic formula

OpenStudy (anonymous):

thank you perl :)